AO Commissioning on 12 October 2011

Observer(s):C. Arcidiacono, L. Busoni, A. Puglisi, JC. Guerra, K. Boutsia, J. Brynnel (LBTO)

Telescope Operator: D. Huerta (LBTO)

SW Support: T. Sargent (Tucson)

Telescope Support: M. Midkiff (LBTO)

Instrument Support: J. Morris (LBTO)

Summary

The sky is clear during sunset. The seeing has been variable, between 0.6" and 1.7"
Pisces is cold but we have to re-install the mechanincal encoder at the filter wheel. We go up the dome at 6:30PM to install the encoder.
We had problems homimg the filter wheel using the pisces interface. We contacted Craig who stayed on the phone while we manually identified the filter position and we set it to FeII.
The filter wheel is currently not connected to the power and it will stay like this for tonight.
At the beginning of the night it took 1h20min for pointing and collimation.
During the night we tested the observe routines, offsets, measured some strehls, tested how to keep the loop during ADI, corrected autogain for faint stars and tested routines with the dither commands. We have always used the filter FeII+ND.

Preparations

Chris installed a patch for IIF to fix the logic problems with the RunAO command (BP13).

Tom installed a patch for MCS and MCSGUI to add some additional rotator reset commands and to fix a telemetry bug (BP13).

John H. updated the global offsets for DXM2IRTCF collimation models with new XY translations that make the pointing of the collimation model match the mount pointing model. The new numbers of X=1.2 Y=1.0 were the average of what we saw over the last three nights. (I don't understand why the hub position appears to have moved by ~2 mm since spring.)

Details

03:23 Finished with the filter wheel for tonight - telescope handed to T.O.

03:48 Open Dome

04:00 Pointing Correction - no star found
  • repeat

05:20 Collimated

05:21 preset - failed - Because the TCS is not getting the preset
  • need to do iif_register as soon as idl is started

05:30 send preset again because the previous one failed
  • observe, AO048
  • guide star not found- because probe could not move to position - guide star out of range
  • change guide star
  • send preset again - loop closed
  • problem with the header - it does not copy some of the original strings - FILTER, UTSTART, OBSMODE

05:44 dimm - 0.9''

05:47 trying offsets
  • it took some time to realise that offsets are ABSOLUTE in the observe routine

06:05 shell ripped during a 5'' offset.
  • repeat preset AO48

06:20 header problem solved - repeat preset and go on with offset testing
  • repeat 10 offsets for offset statistics...
  • the pixelated grid of pisces has not been aligned properly to the telescope reference system - a test to determine the exact angle might be useful
  • data is saved in directory /20111012/AO48

06*58 preset canceled- arbitrator crashed while we were trying a 70'' preset
  • repeat preset - failed
  • arbitrator interface -generic error : command time out (code -13107) - error internal to the AO system

07:10 there seems to be an accumulating tilt which is not recovered from offload: reason unknown ... for now...

07:24 go on with the preset test
  • repeat preset AO48
  • preset failed because star did not converge

07:34 repeat preset AO48 * focus test

07:46 -doing a 30'' offset - to get a sky - file is in directory /20111012/AO48_sky

07:48 save a flat of the secondary

07:53 doing the ADI test.
  • 1st try: preset - autocollimate - runao - pause - stop rotator - resume OB -> doesn't work
  • 2nd try: preset - autocollimate - pause - stop rotator - resume OB - runao -> doesn't work
  • 3rd try: preset - autocollimate - pause - stopr rotator - manually close loop by sending preset again from AOS GUI - resume OB - go on with images -> WORKED
  • there is no need to runao from OB - it is done manually from AOS GUI.
  • script that worked:
preset,OBJECT=AO48,RS=0,GS=13,MODE=ace,PA=-179
auto_collimate
pause
camera,EXP=1.0,FRAMES=1,COADDS=1,FILTER=8
getimage,IMAGE=SCIENCE
  • go on taking images with the rotator off to see if there is a drift in the image - if there is a drift it is NOT caused by the rotator.
  • 90 images - total elapsed time

08:38 changing target - AO159
  • image seems to be saturated - we probably need to change star.

09:09 moving to a faint star (11.4mag)- AO904
  • preset failed - timeout - arbtrator interface generic error - AO did not close the loop in the expected time...
  • repeat preset - error: active optics loop not running - it doesn't find the guide star.
  • change guide star - repeat preset
  • problems with the offload of the tilt. - the offload is not the reason.... trying to find out what is causing the problem - call A.Riccardi
  • while trying to get_image - the loop opened because of "skip frame"

10:26 - trying to solve the problem, we get an image which does not look very good - very bad PSF
  • when we offset to take a sky frame - we got an error - the system was on hold because of the previous time out.

10:44 continue to do preset and autogain on the faint star, in order to see why it fails so often.
  • DIMM - 1.5" seeing

11:07 error with the exp times in the last batch of images... too long - 101sec/8frames instead of 10sec/1frame
  • potential bug: this sequence could not be interrupted, because we would find the pisces server busy... we had to wait for 800sec+overhead
  • these last images were save outside the directory A0904...
  • to take the sky frame we run script A0904_sky. It saves the sky image in there... the next image was saved both in A0904_sky and in the UTdir
  • need to understand at what point the separate commands start saving images outside the directory *.obs.

11:25 The new autogain has been loaded
  • repeat preset to AO904
  • the PSF has improved

11:37 Change star - Astar_29 (8.2mag)
  • preset and take strehl measurements
  • autogain
  • observe, saves the images twice... one in the main directory /PISCES_Data/UTdate and one in the directory with the *obs name. - inform Doug!

12:16 testing dither patterns with the observe routines.
  • it seems to be working - we need to understand how the initial point is set
  • interrupt testing because it is getting too bright

12:53 closing dome

-- KonstantinaBoutsia - 12 Oct 2011
Topic revision: r7 - 15 Oct 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback